# Bridging Semantic Gaps for Language-Supervised Semantic Segmentation

### v1

@@ Abstract
Vision-Language Pre-training has demonstrated its remarkable zero-shot recognition ability and potential to learn generalizable visual representations from language supervision. Taking a step ahead, language-supervised semantic segmentation enables spatial localization of textual inputs by learning pixel grouping solely from image-text pairs. Nevertheless, the state-of-the-art suffers from clear semantic gaps between visual and textual modality: plenty of visual concepts appeared in images are missing in their paired captions. Such semantic misalignment circulates in pre-training, leading to inferior zero-shot performance in dense predictions due to insufficient visual concepts captured in textual representations. To close such semantic gap, we propose Concept Curation (CoCu), a pipeline that leverages CLIP to compensate for the missing semantics. For each image-text pair, we establish a concept archive that maintains potential visually-matched concepts with our proposed vision-driven expansion and text-to-vision-guided ranking. Relevant concepts can thus be identified via cluster-guided sampling and fed into pre-training, thereby bridging the gap between visual and textual semantics. Extensive experiments over a broad suite of 8 segmentation benchmarks show that CoCu achieves superb zero-shot transfer performance and greatly boosts language-supervised segmentation baseline by a large margin, suggesting the value of bridging semantic gap in pre-training data. Code will be made available.
@@ 1 Introduction
We propose Concept Curation (CoCu), a novel pipeline that side-steps the negative effect by semantic bias while exploiting vision-language foundation models for semantic segmentation. CoCu consists of three sequential stages: 1) vision-driven expansion that constructs concept archive via cross-image retrieval; 2) text-to-vision-guided ranking that scores the retrieved concepts according to their assigned relevancies; and 3) cluster-guided sampling that exploits semantic diversity beyond the relevancy scores for concept ranking (Fig. 1 (c)). We perform pre-training from scratch on the segmentation backbone of [43] and evaluate zero-shot transfer over 8 widely adopted segmentation benchmarks. The experiments show that the proposed CoCu improves the baseline as well as the state-of-the-art consistently by large margins, indicating the necessity of closing the semantic gap and the effectiveness of our designs in concept archiving and concept ranking.
@@ 2 Related Works
Vision-Language Pre-training. Recently, Vision-Language Pre-training has become a predominant trend by learning visual representations from natural language supervision [34, 21, 24, 47, 1, 31, 10]. By matching billion-scale image-text pairs via contrast, the learnt representations can be seamlessly transferred to various downstream classification tasks in a zero-shot manner. As a representative, CLIP [34] can match the performance of supervised baselines on ImageNet [13], meanwhile obtain competitive performance over plethora of downstream tasks without accessing any target data. The same learning paradigm has recently been explored for the task of semantic segmentation by hierarchical grouping [43], supervision mining [35, 44, 7, 32], etc. Nevertheless, state-of-the-art language-supervised segmentation is held back by the cross-modal semantic gap between textual and visual pre-training data. Instead of relaxing the strict one-to-one correspondence in vanilla contrastive learning [35], we mitigate the semantic gap by automated curation of relevant visual concepts through concept expanding and concept ranking.

### v4

# Rewrite Caption Semantics: Bridging Semantic Gaps for Language-Supervised Semantic Segmentation

@@ Abstract
Vision-Language Pre-training has demonstrated its remarkable zero-shot recognition ability and potential to learn generalizable visual representations from language supervision. Taking a step ahead, language-supervised semantic segmentation enables spatial localization of textual inputs by learning pixel grouping solely from image-text pairs. Nevertheless, the state-of-the-art suffers from clear semantic gaps between visual and textual modality: plenty of visual concepts appeared in images are missing in their paired captions. Such semantic misalignment circulates in pre-training, leading to inferior zero-shot performance in dense predictions due to insufficient visual concepts captured in textual representations. To close such semantic gap, we propose Concept Curation (CoCu), a pipeline that leverages CLIP to compensate for the missing semantics. For each image-text pair, we establish a concept archive that maintains potential visually-matched concepts with our proposed vision-driven expansion and text-to-vision-guided ranking. Relevant concepts can thus be identified via cluster-guided sampling and fed into pre-training, thereby bridging the gap between visual and textual semantics. Extensive experiments over a broad suite of 8 segmentation benchmarks show that CoCu achieves superb zero-shot transfer performance and greatly boosts language-supervised segmentation baseline by a large margin, suggesting the value of bridging semantic gap in pre-training data. Code is available at https://github.com/xing0047/rewrite.
@@ 1 Introduction
We propose Concept Curation (CoCu), a novel pipeline that side-steps the negative effect by semantic bias111to clarify, we refer to semantic gap as a problem in web-crawled image-text pairs and semantic bias as an issue in pre-trained vision-language model. while exploiting vision-language foundation models for semantic segmentation. CoCu consists of three sequential stages: 1) vision-driven expansion that constructs concept archive via cross-image retrieval; 2) text-to-vision-guided ranking that scores the retrieved concepts according to their assigned relevancies; and 3) cluster-guided sampling that exploits semantic diversity beyond the relevancy scores for concept ranking (Fig. 1 (c)). We perform pre-training from scratch on the segmentation backbone of [43] and evaluate zero-shot transfer over 8 widely adopted segmentation benchmarks. The experiments show that the proposed CoCu improves the baseline as well as the state-of-the-art consistently by large margins, indicating the necessity of closing the semantic gap and the effectiveness of our designs in concept archiving and concept ranking.
@@ 2 Related Works
Vision-Language Pre-training. Recently, Vision-Language Pre-training has become a predominant trend by learning visual representations from natural language supervision [34, 21, 24, 47, 1, 31, 10, zhang2023vision]. By matching billion-scale image-text pairs via contrast, the learnt representations can be seamlessly transferred to various downstream classification tasks in a zero-shot manner. As a representative, CLIP [34] can match the performance of supervised baselines on ImageNet [13], meanwhile obtain competitive performance over plethora of downstream tasks without accessing any target data. The same learning paradigm has recently been explored for the task of semantic segmentation by hierarchical grouping [43], supervision mining [35, 44, 7, 32], etc. Nevertheless, state-of-the-art language-supervised segmentation is held back by the cross-modal semantic gap between textual and visual pre-training data. Instead of relaxing the strict one-to-one correspondence in vanilla contrastive learning [35], we mitigate the semantic gap by automated curation of relevant visual concepts through concept expanding and concept ranking.
@@ 3.3 Concept Curation
To bridge semantic gaps in image-text pairs, we propose Concept Curation (CoCu) to rewrite caption semantics with the help of a pre-trained vision-language model. Consequently, CoCu finds more concept candidates that are aligned to images and compensate for missing semantics in captions. In pre-training, a multi-modal segmentor matches images and visual-enriched captions by contrastive objectives mentioned in Sec. 3.1, encoding better vision-language alignment in its representations. Details of CoCu are described as below.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This project is funded by the Ministry of Education Singapore, under the Tier-2 project scheme with a project number MOE-T2EP20220-0003.